A proton and an electron form two corners of an equilateral triangle with sides of length 3

µm. What is the magnitude of their net electric field at the third corner?


1
Expert's answer
2021-04-26T17:12:23-0400
"E_{net}=2E_x=2\frac{kq_e}{a^2}\cos{60}\\
E_{net}=2\frac{(9\cdot10^{9})(1.6\cdot10^{-19})}{(3\cdot10^{-6})^2}\cos{60}=160\frac{N}{C}"